# Collaboratively adding new knowledge to an LLM

### v1

@@ Abstract
We address the question of how to successively add new knowledge to an LLM whilst retaining previously-added knowledge. We consider two settings, semi-cooperative and fully-cooperative. Overall, LoRA performs better in most cases than full-fine tuning of all parameters when both new knowledge acquisition and retention of old, including recent, knowledge are taken into account. In the semi-cooperative setting, where datasets are not available after training, MOE mixing, model merging, and LoRA-based orthogonal subspace sequential learning, using a small weight on the orthogonality term, perform well. In the fully-cooperative setting where datasets remain available, joint training and sequential training with replay are both effective approaches with LoRA training generally preferable to full fine-tuning. The codes needed to reproduce the results are provided in an open source repository.
@@ 1 Introduction
With the availability of open weight large language models (LLM) such as the popular llama models Touvron et al. (2023), along with easy-to-use libraries for fine-tuning such as Hugging Face Transformers Wolf et al. (2020); Face , training an LLM on one’s own data has become a standard practice. Typically, each use case results in a different fine-tuning of the source model, resulting in many tuned models obtained from the same base model.
LLM tuning aims to give new knowledge or skills to the LLM. The LLM should not, however, lose its ability to perform well on the tasks it was previously trained for, including tasks pertaining to the the general knowledge from pre-training and skills like instruction following. This tug-of-war is further amplified when the LLM is successively subjected to more tuning, over time. While one can return to the base model for each successive tuning, users would like to be able to leverage on recent knowledge from previous tuning, and further tune from there.
@@
The paper proceeds as follows. First we cover the necessary background and related work. Then, in Section 3 we define the methods that we explore. Section 4 presents the experimental setup and the results. Finally, we conclude with a discussion in 5. The codes needed to reproduce the results are provided in an open source repository at https://github.com/RhuiDih/cl.
@@ 2 Related Work
Full fine-tuning a large language model (LLM) involves updating potentially all of its parameters. The seminal work of Hu et al. (2021) introduced LoRA, i.e., Low-Rank Adaptation, a form of parameter-efficient fine-tuning (PEFT), which demonstrated that remarkably good results can be achieved while training only a very small number of parameters associated with the linear layers of the LLM. The success of LoRA paved the way for more efficient training of LLMs.
@@
In parallel, however, the development and widespread use of LoRA also led to questions about whether fine-tuning such a small number of parameters from the base model, instead of all the parameters, would help to reduce the degradation of old knowledge that can arise from training. In Biderman et al. (2024), the authors discuss that question in detail. They conclude that LoRA results in less knowledge acquisition than full fine-tuning (FFT), but that LoRA better retains the base model performance from old knowledge.
@@
The authors of Unsloth , however, claim that the LoRA results in the paper Biderman et al. (2024) were not as good as they could be due to the choice of hyper-parameters used. Specifically, they claim that the implementation of LoRA used in Biderman et al. (2024), which only trained the attention, and up and down matrices, and not the gate projection matrix, performed poorly because the latter is important in obtaining good results. Secondly, they claim that the LoRA training should have included the lm_head and embed_tokens to take into account domain data distribution shift. They state, however, that one must use a smaller learning rate for the lm_head and embed_tokens, and show that training those layers is very sensitive and can result in worse performance in some cases. They also point out that the LoRA rank used in the code tests was $r=256$ with an $\alpha$ of 32, whereas for the math tests, $r=256$ and $\alpha=2r$. They claim that the higher $\alpha$ in the presence of large rank explains the better math results and should also have been used in the code tests. The results of Biderman et al. (2024); Unsloth are important in assessing the benefits and disadvantages of full fine-tuning versus LoRA PEFT. However, we are concerned with how to successively add new knowledge to a model while benefiting from the knowledge added already by earlier fine-tuning. In this context either FFT or PEFT can be used.
@@
A suite of datasets for training and testing was provided in Wang et al. (2023b). The authors analyse model performance across numerous models looking primarily at sequential FFT, sequential LoRA training, and sequential FFT with replay. However, the results provided in the paper are mixed and somewhat conflicting.
@@
Finally, we mention a related work on LoRA training for successive tasks which seeks to drive the adapter to be orthogonal to previously-trained adapters, called o-lora Wang et al. (2023a). We examine whether this technique is of value in the setting where a user does not have access to datasets used in previous training.
@@ 3 Methods
In the semi-collaborative setting, datasets from previous training are not available. Sequential training, whether FFT or LoRA, can be performed. Orthogonal-subspace o-lora Wang et al. (2023a) is a feasible method as well, whereby successive LoRA adapters are encouraged to be orthogonal to preceding adapters by modifying the loss function to be:
@@
such that $O_{i,t}=A_{i}^{T}A_{t}=0$ where $O_{i,t}[j,k]$ denotes the element at the $j$th row and $k$th column of $O_{i,t}$. This method aims to reduce interference between successive LoRA training, postulated to be useful when the tasks of each training are quite different from each other. Note that the orthogonality term $L_{orth}$ in the loss function acts as a soft constraint, and modifies the optimization of the CE loss. As such, we expect that the learning ability of the model trained using o-lora will be reduced in order to achieve the reduced interference with previously-trained adapters.
@@
In addition, two other options are available which do not require access to the datasets used in training the previous models: namely, model merging and MOE model mixing. Model merging refers to averaging the parameters of the model, and a set of methods that extends simple averaging to handle various cases including parameters with opposite signs. The reader may refer to the survey paper Yang et al. (2024) on model merging for more details. In this work, we consider simple parameter averaging as well as the merging method TIES and a combination of TIES Yadav et al. (2023) and DARE Yu et al. (2024a).
@@
MOE model mixing is yet a different method to combine already-trained models without requiring access to the source data used for training them. Model mixing takes one trained model as the base, and several other trained models as experts. The experts contribute the FFN layers (and potentially also the attention layers) in a parallel manner and routers are added to route tokens to a subset of the expert modules. The architecture is the same as that of Mixtral Jiang et al. (2024) without requiring training of the expert modules or the base. Similarly, a Mixture of Experts (MOE) can consist of a set of trained LoRA adapters with a router used to route tokens to a subset of the expert LoRA modules. The authors of Lee et al. (2024) propose several variants for MOE model mixing, some of which require datasets for training the routers, and others which do not, and enable both FFT and LoRA versions of the MOE. We also consider both FFT and LoRA MOE here. Note that the datasets for training the routers need not be the same as those used in training the models. We use the open source repository https://github.ibm.com/Rhui-Dih-Lee/moetify for both the FFT and LoRA MOE.
A distinction to note between the FFT and LoRA MOE is that the FFT MOE becomes large as the number of experts increases. Although inference time remains nearly constant as the number of experts increases since a "top-K" routing policy is generally used, the memory required to store the model increases with number of experts. In our experiments, for this reason, we are limited to a FFT MOE of 4 (out of the possible 7) experts due to memory requirements of including more expert modules, where as the LoRA MOE uses all of the expert modules.
@@
In the fully-collaborative setting, all of the datasets used previously for training the model remain available. In this setting, we consider joint training as an alternative to sequential training. Finally for the fully-collaborative setting, we consider sequential FFT and sequential LoRA training with replay in addition to standard sequential training. Replay is defined as adding a percentage, in our case, 10%, of all previous task data into the current training. Note then that the dataset grows with each successive training. Note also that orthogonal subspace learning, o-lora, is no longer appropriate since previous tasks’ dataset are used in joint and sequential with replay training.
@@ 4.1 Experimental Setup
We make use of the TRACE Wang et al. (2023b) data as part of our experimental setup. The TRACE repository includes 8 tasks, each of which has a 5K-example dataset for training and a 2K-example dataset for testing. The tasks include ScienceQA, which comprises elementary and high school science question-answer pairs, FOMC, a financial hawkish-dovish classification task, MeetingBank, is a city council meeting summarization task, and Py150 a python code-completion task. It also includes two multi-lingual tasks, namely cstance, a Chinese-language zero-shot stance detection task from Chinese social media postings, and 20Minuten, a German news text simplification task.
The TRACE dataset also includes two math tasks that we do not use, making use instead of MetaMath Yu et al. (2024b), a much larger 395K-example dataset for math problem solving, and GSM8K Cobbe et al. (2021) for testing math solving-ability. Hence we consider in total 7 tasks, 6 from TRACE and a math task using a much larger dataset than those in TRACE.
We also make use of evaluation tests from the LM Eval Harness Gao et al. (2024), which includes tests such as MMLU, that reflect the ability of the model to perform well on old knowledge, including the knowledge from pre-training. Experiments are performed on one node with 8 A100-80GB GPU. We use llama3 (meta-llama/Meta-Llama-3-8B) Meta for all tests reported here. The experiments can be reproduced using the code provided in this repository.
@@ 4.2 One-Round Model Training
On the other hand, as shown in Figure 2, old knowledge retention suffers considerably when the training modifies potentially all parameters using full fine-tuning. We take each of the one-round-trained models and assess old knowledge retention through the LM eval harness Gao et al. (2024) suite. Then, we average the results on a per-test basis over the one-round-trained models. We see old knowledge degradation notably on the tasks MMLU and CommonsenseQA which degrade more after FFT than after LoRA training. The final set of bars provides an average over the tests of the model-average scores.
Figure 1: One-round training, starting from the base model, and fine-tuning the model on 6 different dataset-tasks. New knowledge acquisition on average is similar with LoRA and FFT.
Figure 2: Average over the one-round trained models, which were trained from the base model, and each tuned on one dataset-tasks. Average old-knowledge retention after one-round training across models, using full fine-tuning (FFT) suffers considerably more degradation then with LoRA training.
@@ 4.3 Semi-Collaborative Model Training
It is clear that the order of training will have an impact on the resulting performance at any point in the chain of successive training. However, in practice, one may not be able to choose which checkpoint to start from and even if one could choose, the number of choices would become unwieldy as the number of successive training increases. For this reason, we do not explore the impact of task-order on the quality of the output. Instead, we maintain the same order as the authors of the TRACE dataset Wang et al. (2023b) and keep the order constant. The order of training is thus cstance, fomc, meeting, py150, science, 20minuten, and finally the very large math training dataset MetaMath.
We first examine the quality of old knowledge retention with successive training over the 7 datasets and tasks. Figure 3 summarises the results after all 7 rounds of training are complete. As we saw in one-round training, FFT, after successive training, experiences clearly more degradation of old knowledge than LoRA training.
Figure 3: Sequential training: old-knowledge retention after 7 rounds of sequential training, each training starting from the previously-trained model. Full fine-tuning suffers considerably more degradation of old knowledge than LoRA training.
@@
Detailed performance on old knowledge is provided in Table 1 for full fine-tuning and Table 2 for LoRA fine tuning, and for new knowledge acquisition and retention in Table 3 for FFT and Table 4 for LoRA. Observe that overall the degradation from LoRA training is lower, at -1% as compared to -7% with FFT. BWT refers to backward transfer, defined as the percentage change in performance directly after training on the task vs. after all rounds of training have completed.
@@
One way to mitigate this apparent conflict is to add an orthogonality term to the loss function as proposed in Wang et al. (2023a). We consider orthogonal LoRA adapters with a moderate weight of $\lambda=0.5$ on the orthogonality term and a version with a stronger weight of $\lambda=5$ on the orthogonality term.
@@
We also consider four types of model merging here, simple averaging of full model weights, called FFT merge in the figures, analogous averaging of LoRA adapter weights, called LoRA merge, the so-called TIES Yadav et al. (2023) model merging method, and the combination of TIES with another merging method, DARE Yu et al. (2024a), both of which heuristically try to improve the averaging process by dropping some weights or accounting for different signs.
A summary is provided in Figure 7. The TIES and TIES-DARE heuristics do not perform well on new knowledge retention as compared to the other methods. We see that overall, MOE model mixing performs best, both FFT and LoRA MOE Mixing, with the orthogonal subspace o-lora variants performing well also. It should be noted that the FFT MOE Mixed model does not include all 7 of the expert modules, due to the high memory requirement of loading the significantly larger model. Rather we use 4 of the 7 experts in the FFT MOE Mixed model, namely meeting, py150, science and math, hence the average number is not fully performance on new knowledge of the MOE Mix FFT is not fully compaable to the other new knowledge averages. Detailed figures which include task-specific performance are provided in the appendix in Figure 9 for new knowledge acquisition and retention and Figure 10 for old knowledge retention. Full fine-tuning merging via straight averaging performs quite well on old knowledge except for a sharp decline in mmlu which impacts its overall performance.
@@
Figure 7: Semi-cooperative setting. Summary comparison of the final gain from base between orthogonal subspace o-lora successive training versus model merging and MOE model mixing of the individually-trained models, both LoRA MOE and FFT MOE. MOE model mixing overall provides the most benefit by leveraging on synergies across the models.

### v2

@@ Abstract
We address the question of how to successively add new knowledge to an LLM whilst retaining previously-added knowledge. We consider two settings, semi-cooperative and fully-cooperative. Overall, LoRA performs better in most cases than full-fine tuning of all parameters when both new knowledge acquisition and old, including recent, knowledge retention are taken into account. In the semi-cooperative setting, where datasets are not available after training, MOE mixing, model merging, and LoRA-based orthogonal subspace sequential learning, using a small weight on the orthogonality term, perform well. In the fully-cooperative setting where datasets remain available, joint training and sequential training with replay are both effective approaches with LoRA training generally preferable to full fine-tuning.
@@ 1 Introduction
With the availability of open weight large language models (LLM) such as the popular llama models [1], along with easy-to-use libraries for fine-tuning such as Hugging Face Transformers [2, 3], training an LLM on one’s own data has become a standard practice. Typically, each use case results in a different fine-tuning of the source model, resulting in many tuned models obtained from the same base model.
LLM tuning aims to give new knowledge or skills to the LLM. The LLM should not, however, lose its ability to perform well on the tasks it was previously trained for, including tasks pertaining to the general knowledge from pre-training and skills like instruction following. This tug-of-war is further amplified when the LLM is successively subjected to more tuning, over time. While one can return to the base model for each successive tuning, users would like to be able to leverage on recent knowledge from previous tuning, and further tune from there.
@@
The paper proceeds as follows. First we cover the necessary background and related work. Then, in Section 3 we define the methods that we explore. Section 4 presents the experimental setup and the results. Finally, we conclude with a discussion in 5.
@@ 2 Related Work
Full fine-tuning a large language model (LLM) involves updating potentially all of its parameters. The seminal work of [4] introduced LoRA, i.e., Low-Rank Adaptation, a form of parameter-efficient fine-tuning (PEFT), which demonstrated that remarkably good results can be achieved while training only a very small number of parameters associated with the linear layers of the LLM. The success of LoRA paved the way for more efficient training of LLMs.
@@
In parallel, however, the development and widespread use of LoRA also led to questions about whether fine-tuning such a small number of parameters from the base model, instead of all the parameters, would help to reduce the degradation of old knowledge that can arise from training. In [5], the authors discuss that question in detail. They conclude that LoRA results in less knowledge acquisition than full fine-tuning (FFT), but that LoRA better retains the base model performance from old knowledge.
@@
The authors of [6], however, claim that the LoRA results in the paper [5] were not as good as they could be due to the choice of hyper-parameters used. Specifically, they claim that the implementation of LoRA used in [5], which only trained the attention, and up and down matrices, and not the gate projection matrix, performed poorly because the latter is important in obtaining good results. Secondly, they claim that the LoRA training should have included the lm_head and embed_tokens to take into account domain data distribution shift. They state, however, that one must use a smaller learning rate for the lm_head and embed_tokens, and show that training those layers is very sensitive and can result in worse performance in some cases. They also point out that the LoRA rank used in the code tests was $r=256$ with an $\alpha$ of 32, whereas for the math tests, $r=256$ and $\alpha=2r$. They claim that the higher $\alpha$ in the presence of large rank explains the better math results and should also have been used in the code tests. The results of [5, 6] are important in assessing the benefits and disadvantages of full fine-tuning versus LoRA PEFT. However, we are concerned with how to successively add new knowledge to a model while benefiting from the knowledge added already by earlier fine-tuning. In this context either FFT or PEFT can be used.
@@
The authors of [7] analyse model performance across numerous models looking primarily at sequential FFT, sequential LoRA training, and sequential FFT with replay. However, the results provided in the paper are mixed and somewhat conflicting. They provide a suite of datasets for training and testing that we make use of in this work.
@@
Finally, we mention a related work on LoRA training for successive tasks which seeks to drive the adapter to be orthogonal to previously-trained adapters, called o-lora [8]. We examine whether this technique is of value in the setting where a user does not have access to datasets used in previous training.
@@ 3 Methods
In the semi-collaborative setting, datasets from previous training are not available. Sequential training, whether FFT or LoRA, can be performed. Orthogonal-subspace o-lora [8] is a feasible method as well, whereby successive LoRA adapters are encouraged to be orthogonal to preceding adapters by modifying the loss function to be:
@@
The orthogonality term thus aims to enforce $O_{i,t}=A_{i}^{T}A_{t}=0$ to an extent depending upon the parameter $\lambda$, where $O_{i,t}[j,k]$ denotes the element at the $j$th row and $k$th column of $O_{i,t}$. This method aims to reduce interference between successive LoRA training, postulated to be useful when the tasks of each training are quite different from each other. Note that the orthogonality term $L_{orth}$ in the loss function acts as a soft constraint, and modifies the optimization of the CE loss. As such, we expect that the learning ability of the model trained using o-lora will be reduced in order to achieve the reduced interference with previously-trained adapters.
@@
In addition, two other options are available which do not require access to the datasets used in training the previous models: namely, model merging and MOE model mixing. Model merging refers to averaging the parameters of the model, and a set of methods that extends simple averaging to handle various cases including parameters with opposite signs. The reader may refer to the survey paper [9] on model merging for more details. In this work, we consider simple parameter averaging as well as the merging method TIES and a combination of TIES [10] and DARE [11].
@@
MOE model mixing is yet a different method to combine already-trained models without requiring access to the source data used for training them. Model mixing takes one trained model as the base, and several other trained models as experts. The experts contribute the FFN layers (and potentially also the attention layers) in a parallel manner and routers are added to route tokens to a subset of the expert modules. The architecture is the same as that of Mixtral [12] without requiring training of the expert modules or the base. The authors of [13] propose several variants for MOE model mixing, some of which require datasets for training the routers, and others which do not. Training the routers, if desired, is the same as any continued pre-training or supervised fine-tuning, where only the parameters of the routers are able to be modified. Note that datasets used for training the routers need not be the same as those used in training the models. The authors also enable the same approach to be performed using adapters as experts instead of full-parameter models, and release an open-source repo.
@@
In the fully-collaborative setting, all of the datasets used previously for training the model remain available. In this setting, we consider joint training as an alternative to sequential training. Finally for the fully-collaborative setting, we consider sequential FFT and sequential LoRA training with replay in addition to standard sequential training. Replay is defined as adding a percentage, in our case, 10%, of all previous task data into the current training. Note then that the dataset grows with each successive training. Note also that orthogonal subspace learning, o-lora, is no longer appropriate since previous tasks’ dataset are used in joint and sequential training with replay.
@@ 4.1 Experimental Setup
We make use of the TRACE [7] data as part of our experimental setup. The TRACE repository includes 8 tasks, each of which has a 5K-example dataset for training and a 2K-example dataset for testing. The tasks include ScienceQA, which comprises elementary and high school science question-answer pairs, FOMC, a financial hawkish-dovish classification task, MeetingBank, a city council meeting summarization task, and Py150 a python code-completion task. It also includes two multi-lingual tasks, namely cstance, a Chinese-language zero-shot stance detection task from Chinese social media postings, and 20Minuten, a German news text simplification task.
The TRACE dataset also includes two math tasks that we do not use. Instead we make use of MetaMath [14], a much larger 395K-example dataset, for math problem solving extended pre-training, and the two math solving-ability tests, GSM8K and GSM8K_COT [15]. Hence we consider in total 7 tasks, 6 from TRACE and a math task using a much larger dataset than those in TRACE.
We also make use of evaluation tests from the LM Eval Harness [16], which includes tests such as MMLU, arc-challenge, commonsenseQA, hellaswag, openbookQA, piQA, and winogrande, that reflect the ability of the model to perform well on old knowledge, including the knowledge from pre-training. We call these tests collectively "Old Knowledge". Experiments are performed on one node with 8 A100-80GB GPU. We use llama3 (meta-llama/Meta-Llama-3-8B) [17] for all tests reported here.
@@ 4.2 One-Round Model Training
On the other hand, as shown in Figure 2, old knowledge retention suffers considerably when the training modifies potentially all parameters using full fine-tuning. We take each of the one-round-trained models and assess old knowledge retention through the LM eval harness [16] suite. Then, we average the results on a per-test basis over the one-round-trained models. We see old knowledge degradation notably on the tasks MMLU and CommonsenseQA which degrade more after FFT than after LoRA training. The final set of bars provides an average over the tests of the model-average scores.
Figure 1: One-round training, comparing the base model, called "base" in the figure, with performance after fine-tuning the model on 6 different dataset-tasks. New knowledge acquisition on average is similar with LoRA and FFT.
Figure 2: Average over the one-round trained models, which were trained from the base model, and each tuned on one dataset-tasks. Average old-knowledge retention after one-round training across models, using full fine-tuning (FFT) suffers considerably more degradation then with LoRA training. Performance of the original base model is shown in the figure as "base".
@@ 4.3 Semi-Collaborative Model Training
It is clear that the order of training will have an impact on the resulting performance at any point in the chain of successive training. However, in practice, one may not be able to choose which checkpoint to start from and even if one could choose, the number of choices would become unwieldy as the number of successive training increases. For this reason, we do not explore the impact of task-order on the quality of the output. Instead, we maintain the same order as the authors of the TRACE dataset [7] and keep the order constant. The order of training is thus cstance, fomc, meeting, py150, science, 20minuten, and finally the very large math training dataset MetaMath.
We first examine the quality of old knowledge retention with successive training over the 7 datasets and tasks and compare against the scores obtained by the base model, that is, the original llama3.1 model without further training. Figure 3 summarises the results after all 7 rounds of training are complete. As we saw in one-round training, FFT, after successive training, clearly experiences more degradation of old knowledge than LoRA training.
Figure 3: Sequential training: old-knowledge retention after 7 rounds of sequential training, each training starting from the previously-trained model. Full fine-tuning suffers considerably more degradation of old knowledge than LoRA training. "Base" refers to the scores obtained by the original llama3.1 model without further training.
@@
Detailed performance on old knowledge is provided in Table 1 for full fine-tuning and Table 2 for LoRA fine tuning, and for new knowledge acquisition and retention in Table 3 for FFT and Table 4 for LoRA. Observe that overall the degradation from LoRA training is lower, at -1% as compared to -7% with FFT. BWT refers to backward transfer, defined as the percentage change in performance directly after training on the task vs. the rounds of training after the task training. Note that as BWT shows the change due to subsequent training after a task’s own training, BWT is not defined for the last task, which is math, as no training is performed after math.
@@
One way to mitigate this apparent conflict is to add an orthogonality term to the loss function as proposed in [8]. We consider orthogonal LoRA adapters with a moderate weight of $\lambda=0.5$ on the orthogonality term and a version with a stronger weight of $\lambda=5$ on the orthogonality term.
@@
We also consider four types of model merging here, simple averaging of full model weights, called FFT merge in the figures, analogous averaging of LoRA adapter weights, called LoRA merge, the so-called TIES [10] model merging method, and the combination of TIES with another merging method, DARE [11], both of which heuristically try to improve the averaging process by dropping some weights or accounting for different signs.
A summary is provided in Figure 7. The TIES and TIES-DARE heuristics do not perform well on new knowledge retention as compared to the other methods. We see that overall, MOE model mixing performs best, both FFT and LoRA MOE Mixing, with the orthogonal subspace o-lora variants performing well also. It should be noted that the FFT MOE Mixed model does not include all 7 of the expert modules, due to the high memory requirement of loading the significantly larger model. Rather we use 4 of the 7 experts in the FFT MOE Mixed model, namely meeting, py150, science and math, hence the average number is not fully performance on new knowledge of the MOE Mix FFT is not fully comparable to the other new knowledge averages, and care should be taken in assessing MOE Mix FFT with respect to the other methods.
Old Knowledge change is small but positive, reflecting improvement in Old Knowledge performance, when using all of the methods except o-lora with a large weight of 5 and LoRA-DARE-TIES. For those two methods there is a small decrease in Old Knowledge performance.
Detailed figures are provided in the appendix in Figure 9 for new knowledge acquisition and retention and Figure 10 for old knowledge retention. Full fine-tuning merging via straight averaging performs quite well on old knowledge except for a sharp decline in mmlu which impacts its overall performance.
@@
Figure 7: Semi-cooperative setting. Summary comparison of the final gain from base between orthogonal subspace o-lora successive training versus model merging and MOE model mixing of the individually-trained models, both LoRA MOE and FFT MOE. MOE model mixing overall provides the most benefit by leveraging on synergies across the models. Old Knowledge sees a small improvement using most of the methods, with a small decline only using o-lora-5 and LoRA-DARE-TIES.